# Model-independent unbinned analysis of $B\to K^{*}(\to K^{+}\pi^{-})\mu^{+}\mu^{-}$: zeroes, bounds, Wilson coefficients and symmetries

### v1

@@ IV.2 Zero-crossing points of the optimised observables
and using that $T_{1}(q^{2})/V(q^{2})=m_{B}/(m_{B}+m_{K^{*}})$ in this same limit, all form factor dependence drops out and one gets,
@@
This is true up to corrections of ${\cal O}(P_{i}^{CP},P_{3},P_{6,8}^{\prime})$, i.e., the same approximation as in Eq. 55.
@@
${q_{0}^{2}}\Big{\rvert}_{P_{4}^{\prime}}=-2m_{b}{\cal C}_{7}^{\rm eff}\frac{(2m_{b}{\cal C}_{7}^{\rm eff}+m_{B}{\rm Re}{\cal C}_{9}^{\rm eff}(q_{0}^{2}))}{({{\cal C}_{10}^{\rm eff}}^{2}+{|{\cal C}_{9}^{\rm eff}(q_{0}^{2})|}^{2})}$ (68)
in excellent agreement with the next-to-leading order SM prediction, as can be seen in Fig. 13.
@@ Appendix A SM UNBINNED PREDICTIONS OF NON-OPTIMISED OBSERVABLES
For completeness, we provide also the SM predictions of the unbinned non-optimised observables.
@@ Appendix D DISENTANGLING S-WAVE CONTRIBUTIONS AND MASSIVE LEPTONS
To demonstrate the treatment of data with an S-wave component, the pseudo-experiments of the main text have been augmented with a simple S-wave model. The S-wave contribution relative to the P-wave is set to be 0 at the kinematic endpoints ($q^{2}=0\,\text{GeV}^{2}$ and $q^{2}=19\,\text{GeV}^{2}$) and $\sim 10\%$ at $q^{2}=9\,\text{GeV}^{2}$. Here, the interference between P- and S-wave has been neglected in the generation of the pseudo-data for the sake of simplicity. The P-wave events are generated with a relativistic Breit Wigner in $m_{K\pi}$ and the S-wave with a LASS parametrisation.
In the case that lepton masses are non-negligible one must extract $S_{1}^{c}$, $S_{1}^{s}$, $S_{2}^{c}$ and $S_{2}^{s}$ as they are not trivially related to each other. The respective angular functions of these observables are not orthogonal, such that they can only be obtained from a combination of the moments. In the absence of S-wave contributions, the angular observables are obtained from the moments as per Eq. III.
@@
By taking the leptons as massless one can study the $m_{K\pi}$ dependence of the observables for $q^{2}>\sim 4\,\text{GeV}^{2}$, which achieves several aims. The $m_{K\pi}$ dependence of the P-wave observables will be assessed for the first time, shedding new light on the narrow-width approximation for the $K^{\ast{}0}$. One can also extract new information about the S-wave observables, in particular if the commonly used functions (the LASS parametrisation or some isobar model) are accurate. It will also become clear if the amplitudes that describe the S-wave should be considered to be the same at low and high values of $m_{K\pi}$, as is currently assumed by the factorisation of $m_{K\pi}$ from the other kinematic variables in the calculation of the angular observables. The main text highlights the possibility of achieving a fully two-dimensional map in $m_{K\pi}$ and $q^{2}$ of the P-wave observables (without assuming factorisation) in Fig. 6.
@@
Setting $C=0$, one recovers the P-wave only relations of Eq. III. This discussion is moot for the observables describing the interference between the P- and S-wave contributions, $\tilde{S}_{1b}^{c}$, $\tilde{S}_{2b}^{c}$, $\tilde{S}_{4}$, $\tilde{S}_{5}$, $\tilde{S}_{7}$ and $\tilde{S}_{8}$. These are already normalised by $\Gamma_{P}+\Gamma_{S}$ and are orthogonal to the other angular functions and so may be extracted directly from their respective moments.
Using the approach of including $m_{K\pi}$ in the moments calculations, as in Eq. 78 one has the correct extraction of $\tilde{S}_{1a}^{c}$ and $\tilde{S}_{2a}^{c}$, as shown in Fig. 16. Consequently one corrects the normalisation of the P-wave observables with a factor of $1/(1-2\tilde{S}_{1a}^{c}+\frac{2}{3}\tilde{S}_{2a}^{c})$. The effect of this renormalisation, and consequently its importance, is evident in Fig. 17.

### v2

@@ IV.2 Zero-crossing points of the optimised observables
and using that $T_{1}(q^{2})/V(q^{2})=m_{B}/(m_{B}+m_{K^{*}})$ (at leading order in $\alpha_{s}$) in this same limit, all form factor dependence drops out and one gets,
@@
This is true up to corrections of ${\cal O}(P_{i}^{\rm CP},P_{3},P_{6,8}^{\prime})$, i.e., the same approximation as in Eq. (55).
@@
[…]
in excellent agreement with the next-to-leading order SM prediction, as can be seen in Fig. 13. (See, for instance, Ref. Kumar and Mahajan (2016) for explicit leading order expressions of the zero-crossing points of other optimised observables.)
@@ Appendix A SM UNBINNED PREDICTIONS OF NON-OPTIMISED OBSERVABLES
For completeness, we provide in Fig. 14 also the SM predictions of the unbinned non-optimised observables.
@@ Appendix D DISENTANGLING S-WAVE CONTRIBUTIONS AND MASSIVE LEPTONS
To demonstrate the treatment of data with an S-wave component, the pseudo-experiments of the main text have been augmented with a simple S-wave model. The S-wave contribution relative to the P-wave is set to be 0 at the kinematic endpoints ($q^{2}=0\,\text{GeV}^{2}$ and $q^{2}=19\,\text{GeV}^{2}$) and $\sim 10\%$ at $q^{2}=9\,\text{GeV}^{2}$. Here, the interference between P- and S-wave has been neglected in the generation of the pseudo-data for the sake of simplicity. The P-wave events are generated with a relativistic Breit Wigner in $m_{K\pi}$ and the S-wave with a LASS parametrisation Aston et al. (1988); Rui and Wang (2018).
In the case that lepton masses are non-negligible one must extract $S_{1}^{c}$, $S_{1}^{s}$, $S_{2}^{c}$ and $S_{2}^{s}$ as they are not trivially related to each other. The respective angular functions of these observables are not orthogonal, such that they can only be obtained from a combination of the moments. In the absence of S-wave contributions, the angular observables are obtained from the moments as per Eq. (III).
@@
By taking the leptons as massless one can study the $m_{K\pi}$ dependence of the observables for $q^{2}\gtrsim 4\,\text{GeV}^{2}$, which achieves several aims. The $m_{K\pi}$ dependence of the P-wave observables will be assessed for the first time, shedding new light on the narrow-width approximation for the $K^{\ast{}0}$. One can also extract new information about the S-wave observables, in particular if the commonly used functions (the LASS parametrisation or some isobar model) are accurate. It will also become clear if the amplitudes that describe the S-wave should be considered to be the same at low and high values of $m_{K\pi}$, as is currently assumed by the factorisation of $m_{K\pi}$ from the other kinematic variables in the calculation of the angular observables. The main text highlights the possibility of achieving a fully two-dimensional map in $m_{K\pi}$ and $q^{2}$ of the P-wave observables (without assuming factorisation) in Fig. 6.
@@
Setting $C=0$, one recovers the P-wave only relations of Eq. (III). This discussion is moot for the observables describing the interference between the P- and S-wave contributions, $\tilde{S}_{1b}^{c}$, $\tilde{S}_{2b}^{c}$, $\tilde{S}_{4}$, $\tilde{S}_{5}$, $\tilde{S}_{7}$ and $\tilde{S}_{8}$. These are already normalised by $\Gamma_{P}+\Gamma_{S}$ and are orthogonal to the other angular functions and so may be extracted directly from their respective moments.
Using the approach of including $m_{K\pi}$ in the moments calculations, as in Eq. (78) one has the correct extraction of $\tilde{S}_{1a}^{c}$ and $\tilde{S}_{2a}^{c}$, as shown in Fig. 16. Consequently one corrects the normalisation of the P-wave observables with a factor of $1/(1-2\tilde{S}_{1a}^{c}+\frac{2}{3}\tilde{S}_{2a}^{c})$. The effect of this renormalisation, and consequently its importance, is evident in Fig. 17.